The sides of a triangle are 6, 8 and 15 cm long. In a simikar triangle, the longest side is 21 cm long. Find the other two sides.
Answer by Hawksfan(61)

This is a ratio problem. lets call the unknown sides x and y.
Since the 21 is the longest side of 2nd triangle, we compare it to 15, the longest side of the first triangle.
21/15 = x/6
simplify
7/5 = x/6
multiply by 6
x = 42/5

now find the other side y
21/15 = y/8
7/5 = y/8
56/8 =y or y = 56/8 = 7